OpenStudy (anonymous):

Partial Fraction Decomposition: I have everything right except the last linear. Here's the given: (x+4)/[(x^2)(x^2+4)] Here is the system I got from it: A+C=0 B+D=0 4A=1 4B=4 So when I solved for the first three unknowns, I got A = 1/4 B = 1 C = -1/4 The back of the book says those are all correct, except for D, which gave me -1, but the book says it's x+4. How is that? Shouldn't it be x-1?

OpenStudy (dumbcow):

hmm D = -1 is correct \[\frac{\frac{1}{4}x+1}{x^{2}} - \frac{\frac{1}{4}x+1}{x^{2}+4}\] which simplifies to \[\frac{1}{4}(\frac{x+4}{x^{2}} - \frac{x+4}{x^{2}+4})\]

OpenStudy (anonymous):

Oops! I just figured it out. My answer is right, my book just wrote it differently *facepalm* Instead of \[-\frac{ 1 }{ 12 }x - \frac{ 1 }{ 3 }\] which is my answer, it put it in unfactored form, which is \[-\frac{ 1 }{ 12 }(x+4)\] which is the same thing, unless my tired brain is playing tricks on me...

OpenStudy (anonymous):

Oops again, the above is actually for a different problem (gave me the same trouble though).

OpenStudy (anonymous):

Please correct me if I'm wrong, but (and this is for the correct problem) here's my answer: \[-\frac{ 1 }{ 4 }x - 1\] and here's my book's answer: \[-\frac{ 1 }{ 4 }(x+4)\] They're the same thing, right?

OpenStudy (dumbcow):

yes same

OpenStudy (anonymous):

I'll give you a medal just for the assurance you gave me :) Thank you!